# Almost all graphs are vertex-minor universal

### v1

@@ 1 Background and main results
Motivated by this perspective, Bravyi, Sharma, Szegedy, and de Wolf [5] introduced the notion of $k$-pairable quantum states. An $n$-party state $|\psi\rangle$ is $k$-pairable if, for any choice of $k$ disjoint pairs of parties, there exists a LOCC protocol (consisting of local Clifford operations, local Pauli measurements, and classical communications) that starts with $|\psi\rangle$ and ends in a state such that each of those pairs of parties share an EPR pair. This definition reflects a worst-case notion of network universality: a single fixed resource state must, under LOCC, generate EPR pairs for any choice of $k$ disjoint pairs of parties. Bravyi et al. [5] established upper and lower bounds on the achievable value of $k$ in terms of the number of parties and number of qubits per party.
@@
Vertex-minors capture an important sequence of operations on quantum states: if $H$ can be obtained as a vertex-minor of $G$, then the quantum graph state corresponding to $H$ can be obtained from the graph state corresponding to $G$ through a LOCC protocol. The converse also holds if $H$ has no isolated vertices [16].
@@
We emphasize that we require each $H$ to be obtained exactly on each specified vertex set $U$. In particular, there are ${|V(G)|\choose k}2^{{k\choose 2}}$ total choices of pairs $U$ and $H$. If $G$ is a $k$-vertex-minor universal graph then the quantum graph state corresponding to $G$ is $k$-stabilizer universal. This enables explicit constructions of universal resource states for quantum communication protocols [12, 9].
@@
Specifically regarding the notion of $k$-vertex-minor universal graphs, Claudet et al. [12] showed that there exist $k$-vertex-minor universal graphs with $O(k^{4}\log k)$ vertices, while any $k$-vertex-minor universal graph must have at least $(k-2)^{2}/(2\log_{2}3)$ vertices. Cautrès et al. [9] improved the existence result to match the lower bound up to a constant factor, establishing the existence of $k$-vertex-minor universal graphs on $(2+o(1))k^{2}$ vertices. Their construction is based on a random unbalanced bipartite graph model where one side has $\Theta(k^{2})$ vertices and the other has $\Theta(k\log k)$ vertices, with edges between the two sides included independently with probability $1/2$. This yields a graph with $\Theta(k^{3}\log k)$ edges with high probability, which is less than a constant proportion of possible edges.
@@
For all $c>0$, if $G\sim\mathbb{G}(n,1/2)$ with $n\geq(1+c)\frac{2}{\log_{2}(16/13)}k^{2}$, then $G$ is $k$-pivot-minor universal with probability at least $1-2^{-(1+o(1))ck^{2}}$.
@@ 2 Preliminaries and proof overview
We denote the symmetric difference of sets by $A\mathbin{\triangle}B=(A\setminus B)\cup(B\setminus A)$. For a graph $G$, we use $V(G)$ and $E(G)$ to denote its vertex set and edge set respectively. For $U\subseteq V(G)$, we use $G[U]$ to denote the graph induced on $U$, and GUG-U to denote the graph obtained from deleting $U$. The adjacency matrix of a graph $G$ (over $\mathbb{F}_{2}$) is the symmetric matrix $A\in\mathbb{F}_{2}^{V(G)\times V(G)}$ such that $A_{u,v}=1$ if $uv\in E(G)$ and $A_{u,v}=0$ otherwise. Note that as we assume graphs to be simple, $A$ has 0 along the diagonal. For sets $X,Y$, we define the set $X,Y]\coloneqq\{\{x,y\}\mid x\in X,y\in Y\}$ and let $G[X,Y]$ be the induced bipartite subgraph with edge set $E(G)\cap[X,Y]$.
@@ 2.1 Vertex-minor preliminaries
We emphasize that the sequence $(x_{i})_{i}$ depends only on the subgraph $\hat{G}=G[\hat{V}$], and is independent of the edges between $\hat{V}$ and $V(G)\setminus\hat{V}$. Hence, if $G=\mathbb{G}(n,1/2)$, then by revealing only the edges of $\hat{G}$ we can rewrite any arbitrary sequence of complementations $(v_{i})_{i}$ in the desired form $\circ(x_{j})_{j}$, and then study the statistics of the resulting graph. The proof is deferred to Section 3.3.
@@ 2.2 Probability preliminaries
Since the map $G\mapsto G*v$ is an involution on $\mathcal{G}_{n}$, it preserves the uniform distribution. ∎
@@ 3.1 Random walks on the space of graphs
We are now ready to prove our main result via the union bound over all choices of $U\subseteq V(G)$ and graphs $H$ with vertex set $U$.
Proof of Theorem 1.1.
Let $G\sim\mathbb{G}(n,p)$, where $n=(1+c)\frac{1}{2\log_{2}(4/3)}k^{2}$. Here we allow $c=c(k)$ to be increasing with $k$, but it is lower bounded by some positive constant. By Lemma 3.9, the probability that a given graph $H$ on $k$ given vertices of $V(G)$ is not a vertex-minor of $G$ is at most $2^{2k}(3/4)^{n-k}+2^{-n+k+\binom{k}{2}}(n-k)^{2k}=2^{2k}(3/4)^{(1+o(1))n}$. By the union bound, we have
@@ 4 Universality for pivot-minors and binary matroids
Note that this is strictly stronger than only requiring that graph $(U_{1}\cup U_{2},E_{H})$ is a (labeled) pivot-minor of $(L\cup R,E)$. Moreover, recall that because we preserve vertex labels, graphs may be isomorphic without one being a pivot-minor of the other.
@@ 4.1 Reduction from binary matroids to bipartite pivot-minors
We now discuss the connection between pivot-minors in ordered bipartite graphs and minors of binary matroids, reducing Theorem 4.2 to Theorem 4.1. More detail on this relationship an be found in [37]. We assume familiarity with basic matroid theory, see Oxley’s book [38] for reference. All matroid minors will be labeled (not equivalent up to isomorphism). Let $G=(L,R,E)$ be an ordered bipartite graph. Let $A=\operatorname{Adj}(G)\in\{0,1\}^{R\times L}$ be the bipartite adjacency matrix between parts $L$ and $R$. That is, with $A_{u,v}=1$ if and only if $u\in R$ is adjacent to $v\in L$. To each ordered bipartite graph $G=(L,R,E)$ we associate the binary matroid $M(G)=M(L,R,E)$ on ground set $L\mathbin{\sqcup}R$ represented by the matrix $I_{R}\ A]\in\{0,1\}^{R\times V(G)}$ where $I_{R}$ is the $R\times R$ identity matrix. We then say $G$ is the fundamental graph of $M(G)$ with respect to the basis $R$.
@@
This implies that a matroid $M$ is $k$-minor universal if and only if its fundamental graph $G$ is $k$-pivot-minor universal. This is almost sufficient to reduce Theorem 4.2 to Theorem 4.1, but with one technical issue: the uniform distribution on the set $\mathcal{M}_{r,n}$ of rank-$r$ binary matroids on ground set $[n]$ (henceforth $\mu_{r,n}$) is not equivalent to the distribution $\mathbb{G}(\ell,r,1/2)$ on their fundamental graphs. The remainder of this section will show that these two distributions are actually very close.
@@
To show that the distributions $\mathbb{G}(\ell,r,1/2)$ and $\mu_{r,n}$ are not too far apart, it now suffices to show that $b(M)$ is concentrated in the latter distribution. First, we show that $\mu_{r,n}$ is close to the “random representable matroid” model of Kelly and Oxley [32], and also studied in [21, 1]. For $A\in\mathbb{F}_{2}^{r\times[n]}$, let $M(A)$ be the labeled matroid represented by $A$. Note that $M(A)\in\mathcal{M}_{r,n}$ if and only if $A$ is full rank. Finally, it is easy to show [38, Prop. 6.4.1] that each $M\in\mathcal{M}_{r,n}$ is uniquely representable up to row operations, and thus has exactly $|\mathrm{GL}_{r}(2)|$ representation matrices $A\in\mathbb{F}_{2}^{r\times[n]}$, where by $\mathrm{GL}_{r}(2)$ denotes the general linear group of $r\times r$ invertible matrices over $\mathbb{F}_{2}$. This yields the following:
@@
This relationship is notably false for general $\mathbb{F}_{q}$ representable matroids. It is easy to check [32, Lemma 3.1] that $A\in\mathbb{F}_{2}^{r\times n}$ is full rank with probability $\prod_{i=1}^{r}(1-2^{r-n-i})\geq 1-2^{-(n-r)}$, and therefore the uniform distribution is close to the random representable matroid in total variation distance. Next, we concentrate $b(M)$ for a uniformly random or random representable matroid.
@@
The hypergeometric distribution $\mathrm{Hypergeom}(n,r,r)$ (i.e. sampling without replacement) is dominated in convex order by the binomial $\mathrm{Bin}(r,r/n)$ (implied by Heoffding [30, Theorem 4]). Since $x\mapsto 2^{x}$ is convex, we conclude
@@ 5 Vertex-minor Ramsey and concluding remarks
Let $R_{\mathrm{piv}}(s,t)$ be the minimum $n$ such that any $n$-vertex graph contains an independent set of size $s$ as a pivot-minor or a clique of size $t$ as a pivot-minor.
We note that trivially $R_{\mathrm{vm}}(\min\{s,t-1\})\leq R_{\mathrm{piv}}(s,t)\leq R(s,t)$. Theorem 1.2 supports the following conjecture, which would imply 1.5.
@@
$R_{\mathrm{piv}}(s,t)\leq\mathsf{poly}(s,t)$.
We note that proper pivot-minor closed classes also have the Erdős-Hajnal property [17], so this conjecture holds if we restrict to any proper pivot-minor closed class. We also improve the trivial upper bound for $R_{\mathrm{piv}}(s,t)$.
@@ Acknowledgments.
This work began at the Fall 2025 Atlanta Graduate Combinatorics Workshop at Georgia Tech. We thank Georgia Tech for providing funding for the workshop, Sean Longbrake for helping to organize it, and Elliot Frederickson for his motivating presence there. We also extend our thanks to Xiaoyu He and Rose McCarty for enlightening conversations, and to Rose McCarty for telling us about the problem of whether the uniformly random graph is vertex-minor universal (which was originally posed by Nathan Claudet) and pointing us to relevant sources.

### v2

@@ 1 Background and main results
Motivated by this perspective, Bravyi, Sharma, Szegedy, and de Wolf [5] introduced the notion of $k$-pairable quantum states. An $n$-party state $|\psi\rangle$ is $k$-pairable if, for any choice of $k$ disjoint pairs of parties, there exists a LOCC protocol that starts with $|\psi\rangle$ and ends in a state such that each of those pairs of parties share an EPR pair. This definition reflects a worst-case notion of network universality: a single fixed resource state must, under LOCC, generate EPR pairs for any choice of $k$ disjoint pairs of parties. Bravyi et al. [5] established upper and lower bounds on the achievable value of $k$ in terms of the number of parties and number of qubits per party.
@@
Vertex-minors capture an important sequence of operations on quantum states: if $H$ can be obtained as a vertex-minor of $G$, then the quantum graph state corresponding to $H$ can be obtained from the graph state corresponding to $G$ through just local Clifford operations, local Pauli measurements, and classical communications. The converse also holds if $H$ has no isolated vertices [19].
@@
We emphasize that we require each $H$ to be obtained exactly on each specified vertex set $U$. In particular, there are $\binom{|V(G)|}{k}2^{\binom{k}{2}}$ total choices of pairs $U$ and $H$. If $G$ is a $k$-vertex-minor universal graph then the quantum graph state corresponding to $G$ is $k$-stabilizer universal. This enables explicit constructions of universal resource states for quantum communication protocols [12, 9].
@@
Specifically regarding the notion of $k$-vertex-minor universal graphs, Claudet et al. [12] showed that there exist $k$-vertex-minor universal graphs with $O(k^{4}\log k)$ vertices, while any $k$-vertex-minor universal graph must have at least $(k-2)^{2}/(2\log_{2}3)$ vertices. Cautrès et al. [9] improved the existence result to match the lower bound up to a constant factor, establishing the existence of $k$-vertex-minor universal graphs on $(2+o(1))k^{2}$ vertices. Their construction is based on a random unbalanced bipartite graph $G$ where one side has $\Theta(k^{2})$ vertices and the other has $\Theta(k\log k)$ vertices, with edges between the two sides included independently with probability $1/2$. This yields a graph with $\Theta(k^{3}\log k)$ edges with high probability, which is less than a constant proportion of possible edges. They also give a polynomial time algorithm to find a $k$-vertex graph $H$ as a vertex-minor in $G$ with high probability.
@@
For all $c>0$, if $G\sim\mathbb{G}(n,1/2)$ with $n\geq(1+c)\frac{2}{\log_{2}(16/13)}k^{2}\approx(1+c)6.68k^{2}$, then $G$ is $k$-pivot-minor universal with probability at least $1-2^{-(1+o(1))ck^{2}}$.
@@ 2 Preliminaries and proof overview
We denote the symmetric difference of sets by $A\mathbin{\triangle}B=(A\setminus B)\cup(B\setminus A)$. For a graph $G$, we use $V(G)$ and $E(G)$ to denote its vertex set and edge set respectively. For $U\subseteq V(G)$, we use $G[U]$ to denote the graph induced on $U$, and GUG-U to denote the graph obtained from deleting $U$. The adjacency matrix of a graph $G$ (over $\mathbb{F}_{2}$) is the symmetric matrix $A\in\mathbb{F}_{2}^{V(G)\times V(G)}$ such that $A_{u,v}=1$ if $uv\in E(G)$ and $A_{u,v}=0$ otherwise. Note that as we assume graphs to be simple, $A$ has 0 along the diagonal. For disjoint sets $X,Y$, we define the set $X,Y]\coloneqq\{\{x,y\}\mid x\in X,y\in Y\}$ and let $G[X,Y]$ be the induced bipartite subgraph with edge set $E(G)\cap[X,Y]$.
@@ 2.1 Vertex-minor preliminaries
We emphasize that the sequence $(x_{i})_{i}$ depends only on the subgraph $\hat{G}=G[\hat{V}$], and is independent of the edges between $\hat{V}$ and $V(G)\setminus\hat{V}$. Without this careful tracking of the dependency, this lemma would be implied by known results, (see [14, Proposition 6] or [13, Proposition 7]). In Section 3.3, we give a short proof by induction using Theorem 2.2.
Because of the dependence only on $\hat{G}$, if $G=\mathbb{G}(n,1/2)$, then by revealing only the edges of $\hat{G}$ we can rewrite any arbitrary sequence of complementations $(v_{i})_{i}$ in the desired form $\circ(x_{j})_{j}$, and then study the statistics of the resulting graph.
@@ 2.2 Probability preliminaries
Since the map $G\mapsto G*v$ is an involution on $\mathcal{G}_{n}$, it preserves the uniform distribution. Indeed, for any graph $G^{\prime}\in\mathcal{G}_{n}$, we have $\mathbb{P}[G*v=G^{\prime}]=\mathbb{P}[G=G^{\prime}*v]=\frac{1}{|\mathcal{G}_{n}|}$. ∎
@@ 3.1 Random walks on the space of graphs
We are now ready to prove our main result via the union bound over all choices of $U\subseteq V(G)$ and graphs $H$ with vertex set $U$. We first restate the theorem for convenience.
See 1.1
Proof.
Let $G\sim\mathbb{G}(n,p)$, where $n=(1+c)\frac{1}{2\log_{2}(4/3)}k^{2}$. Here we allow $c=c(k)$ to be increasing with $k$, but it is lower bounded by some positive constant. By Lemma 3.9, the probability that a given graph $H$ on $k$ given vertices of $V(G)$ is not a vertex-minor of $G$ is at most $2^{2k}(3/4)^{n-k}+2^{-n+k+\binom{k}{2}}(n-k)^{2k}=2^{2k}(3/4)^{(1+o(1))n}$.
By the union bound, we have
@@ 4 Universality for pivot-minors and binary matroids
Note that this is strictly stronger than only requiring that graph $(U_{1}\cup U_{2},E_{H})$ is a (labeled) pivot-minor of $(L\cup R,E)$, since we require the partitions of $H$ and $G$ to be aligned.
Moreover, recall that because we preserve vertex labels, graphs may be isomorphic without one being a pivot-minor of the other.
@@ 4.1 Reduction from binary matroids to bipartite pivot-minors
We now discuss the connection between pivot-minors in ordered bipartite graphs and minors of binary matroids, reducing Theorem 4.2 to Theorem 4.1. More details on this relationship can be found in [40]. We assume familiarity with basic matroid theory, see Oxley’s book [41] for reference. All matroid minors will be labeled (not equivalent up to isomorphism). Let $G=(L,R,E)$ be an ordered bipartite graph. Let $A=\operatorname{Adj}(G)\in\{0,1\}^{R\times L}$ be the bipartite adjacency matrix between parts $L$ and $R$. That is, with $A_{u,v}=1$ if and only if $u\in R$ is adjacent to $v\in L$. To each ordered bipartite graph $G=(L,R,E)$ we associate the binary matroid $M(G)=M(L,R,E)$ on ground set $L\mathbin{\sqcup}R$ represented by the matrix $I_{R}\ A]\in\{0,1\}^{R\times V(G)}$ where $I_{R}$ is the $R\times R$ identity matrix. We then say $G$ is the fundamental graph of $M(G)$ with respect to the basis $R$.
@@
This implies that a matroid $M$ is $k$-minor universal if and only if its fundamental graph $G$ is bipartite $k$-pivot-minor universal. This is almost sufficient to reduce Theorem 4.2 to Theorem 4.1, but with one technical issue: the uniform distribution on the set $\mathcal{M}_{r,n}$ of rank-$r$ binary matroids on ground set $[n]$ (henceforth $\mu_{r,n}$) is not equivalent to the distribution $\mathbb{G}(\ell,r,1/2)$ on their fundamental graphs. The remainder of this section will show that these two distributions are actually very close.
@@
To show that the distributions $\mathbb{G}(\ell,r,1/2)$ and $\mu_{r,n}$ are not too far apart, it now suffices to show that $b(M)$ is concentrated in the latter distribution. First, we show that $\mu_{r,n}$ is close to the “random representable matroid” model of Kelly and Oxley [35], and also studied in [24, 1]. For $A\in\mathbb{F}_{2}^{r\times[n]}$, let $M(A)$ be the labeled matroid represented by $A$. Note that $M(A)\in\mathcal{M}_{r,n}$ if and only if $A$ is full rank. Finally, it is easy to show [41, Prop. 6.4.1] that each $M\in\mathcal{M}_{r,n}$ is uniquely representable up to row operations, and thus has exactly $|\mathrm{GL}_{r}(2)|$ representation matrices $A\in\mathbb{F}_{2}^{r\times[n]}$, where by $\mathrm{GL}_{r}(2)$ denotes the general linear group of $r\times r$ invertible matrices over $\mathbb{F}_{2}$. This yields the following:
@@
This relationship is notably false for general $\mathbb{F}_{q}$ representable matroids. It is easy to check [35, Lemma 3.1] that $A\in\mathbb{F}_{2}^{r\times n}$ is full rank with probability $\prod_{i=1}^{r}(1-2^{r-n-i})\geq 1-2^{-(n-r)}$, and therefore the uniform distribution is close to the random representable matroid in total variation distance. Next, we concentrate $b(M)$ for a uniformly random or random representable matroid.
@@
The hypergeometric distribution $\mathrm{Hypergeom}(n,r,r)$ (i.e. sampling without replacement) is dominated in convex order by the binomial $\mathrm{Bin}(r,r/n)$ (implied by Heoffding [33, Theorem 4]). Since $x\mapsto 2^{x}$ is convex, we conclude
@@ 5 Vertex-minor Ramsey and concluding remarks
Let $R_{\mathrm{piv}}(k)$ be the minimum $n$ such that any $n$-vertex graph contains an independent set or clique of size $k$ as a pivot-minor.
Note that unlike the definition of $R_{\mathrm{vm}}(k)$, we need to allow either a clique or an independent set in this definition since $K_{n}$ does not contain any non-trivial independent set as a pivot-minor.
We note that trivially $R_{\mathrm{vm}}(k-1)\leq R_{\mathrm{piv}}(k)\leq R(k)$, the diagonal Ramsey number. Theorem 1.2 supports the following conjecture, which would imply 1.5.
@@
$R_{\mathrm{piv}}(k)\leq\mathsf{poly}(k)$.
We note that proper pivot-minor closed classes also have the Erdős-Hajnal property [20], so this conjecture holds if we restrict to any proper pivot-minor closed class. We also improve the trivial upper bound of $R(k)$ for $R_{\mathrm{piv}}(k)$.
@@ Acknowledgments.
This work began at the Fall 2025 Atlanta Graduate Combinatorics Workshop at Georgia Tech. We thank Georgia Tech for providing funding for the workshop, Sean Longbrake for helping to organize it, and Elliot Frederickson for his motivating presence there. We also extend our thanks to Xiaoyu He and Rose McCarty for enlightening conversations, and to Rose McCarty for telling us about the problem of whether the uniformly random graph is vertex-minor universal (which was originally posed by Nathan Claudet) and pointing us to relevant sources. Finally, we thank Nathan Claudet for helpful comments on an earlier version of this paper.